OpenStudy (anonymous):

Please Help me Given that f(x)=1/x and g(x)=8 x - 4, calculate g o f(x) f o f(x) g o g(x) when i solved g o f(x) I ended up with 8(1/x)-4 and that came to 8x-4...which i got wrong For the other twoo I wasn't sure where to even start. won't they just equal eachother? wouldn't f o f(x) just be 1/x=1/x ? I got that wrong too. Help please.

OpenStudy (unklerhaukus):

\[ f(x)=\frac1x,\qquad g(x)=8 x - 4\] \[g\circ f(x)=g(f(x))=g\left(\frac1x\right)=8\left(\frac1x\right)-4=\frac8x-4\] \[f\circ f(x)=f\left(f(x)\right)=f\left(\frac1x\right)=\frac1{\frac1x}=x\]

OpenStudy (anonymous):

Thank you so much. I see how you did the last one, I think my brain is just trying to make sense of it

OpenStudy (anonymous):

I though that it would be 1/x=1/x for some reason

OpenStudy (unklerhaukus):

what do you get for \(g\circ g(x)\)

OpenStudy (anonymous):

I did the same thing and got it wrong. I just did 8x-4=8x-4 ...but I'm thinking the set up may be more like 8(8x-4)-4 ...is that right? then my answer would be (64x-32)-4 which would be 64x-28? I may have just over complicated that

OpenStudy (anonymous):

for the g o f if the x is in the denominator, then wouldn't your domain be all real numbers except -8?

OpenStudy (anonymous):

probably 0 would make more sense

OpenStudy (unklerhaukus):

\[g\circ g(x)=g(8x-4)=8(8x-4)-4=64x-32-4\] is right so far but -32-4 is not -28, try that bit again

OpenStudy (anonymous):

so 64x-36

OpenStudy (unklerhaukus):

\[\checkmark\]

OpenStudy (anonymous):

Thank you so much for taking the time to explain this to me. I was having a serious meltdown trying to understand all this stuff for an exam in 3 weeks. Your are amazing! Thank you so much!!

OpenStudy (unklerhaukus):

now should we look at the domains?

OpenStudy (anonymous):

oh gawd...i feel stressed already. yes we should

OpenStudy (anonymous):

so for g o f i think the domain would be all real numbers except 0, because it can't zero out right? so it would be (-infinity, 0] ?

OpenStudy (unklerhaukus):

all real numbers except 0, is right but you havent got the interval notation right (-infinity, 0] is only half of the domain

OpenStudy (anonymous):

for f of f I think all real numbers except zero...so (-infinity, 0] or would it be (-infinity, 0) since you don't want it to be 0?

OpenStudy (anonymous):

i don't understand, what do you mean by half of the domain?

OpenStudy (unklerhaukus):

x can be positive too

OpenStudy (anonymous):

so then...umm. (-infinity, 0], [0, infinity)

OpenStudy (anonymous):

or would i use the U since fit can't be zero?

OpenStudy (unklerhaukus):

the domain of f(x)=1/x is (-∞,0)U(0,∞) square brackets mean the boundary term Is included, ( and it should't be in this case )

OpenStudy (anonymous):

that makes sense. so i only include the bracket sign when the actual number is included? but isn't the domain always "what can't equal zero" so then wouldn't everything under the domain really be enclosed in ( )

OpenStudy (anonymous):

what i just wrote confused me. so i'll move on to the domain of g o g now

OpenStudy (anonymous):

so it would be: (-infinity, infinity) ...right?

OpenStudy (unklerhaukus):

Well you have got the right domain for g o g , yes (all real numbers)

OpenStudy (unklerhaukus):

the domain for f o f is a little trickier

OpenStudy (anonymous):

whew. i'm glad something made sense. Thank you so much for your help.

OpenStudy (anonymous):

wouldn't it just be ..all real numbers?

OpenStudy (anonymous):

I want to say x for some reason

OpenStudy (unklerhaukus):

\[f\circ f(x)=\frac1{\frac1x}=x\] x cannot be zero

OpenStudy (anonymous):

I don't understand how you came up with that.

OpenStudy (anonymous):

I know it can't be "x" bc that isn't an actual number, it seems like -(x/1) would make more sense

OpenStudy (unklerhaukus):

this page might explain it better than i can , http://www.regentsprep.org/Regents/math/algtrig/ATP7/domaincomposite.htm

OpenStudy (anonymous):

Thank you! I'll be printing that out about now. well it's past midnight here and I have to drag myself to work in the morning so I need to get off, but thank you so much for the help and for the link to that website. I really appreciate it.

OpenStudy (unklerhaukus):

Khan academy is another good resource when you have time https://www.khanacademy.org/math/algebra/algebra-functions good night and good luck

OpenStudy (anonymous):

Thank you very much!